xy^2dx+zx^2dy-x^2y^2dz= determine the following equation are integral
1
Expert's answer
2021-12-13T05:09:18-0500

"xy^{\mathrm{2}}\ dx\ +\ \ zx^{\mathrm{2}}\ dy\ \ \ -\ \ x^{\mathrm{2}}y^{\mathrm{2}}\ dz\ =0 \\ 
\frac{y^{\mathrm{2}}d\left(x^{\mathrm{2}}\right)}{\mathrm{2}}\ \ +\ zx^{\mathrm{2}}\ dy-\ x^{\mathrm{2}}y^{\mathrm{2}}\ dz\ =0 \\ 
 \\ 
\int{\left(\frac{y^{\mathrm{2}}d\left(x^{\mathrm{2}}\right)}{\mathrm{2}}\right)}{}{}+\ \ \int{zx^{\mathrm{2}}dy{}{}{}{}\ \ -\ \ \ \int{\left(x^{\mathrm{2}}y^{\mathrm{2}}\ dz\ \right)=0}} \\ 
 \\ 
\frac{\left(x^{\mathrm{2}}y^{\mathrm{2}}\right)}{\mathrm{2}}+yzx^{\mathrm{2}}{}-{}{}x^{\mathrm{2}}y^{\mathrm{2}}\ z{}{}{}\ \ =\ \ C"